a rectangular garden 10 feet wide and 12 feet long is surrounded by a rock border of uniform width. the area of the garden and rock border combined is 168 square feet. what is the width of the rock border?
Answer by checkley77(12844)

(10+2x)(12+2x)=168
120+44x+4x^2-168=0
4x^2+44x-48=0
4(x^2+11x-12)=0
4(x+12)(x-1)=0
x-1=0
x=1 ft. ans. for the width of the rock border.
Proof:
(10+2*1)(12+2*1)=168
(10+2)(12+2)=168
12*14=168
168=168